A spacelab circles Earth at an altitude of 380 miles. When an astronaut views the horizon of Earth, the angle θ is 65.8 dgrees. Use this information to estimate the radius of the Earth.






Answer by stanbon(48558)

A spacelab circles Earth at an altitude of 380 miles. When an astronaut views the horizon of Earth, the angle θ is 65.8 dgrees. Use this information to estimate the radius of the Earth.
--------------------------------------
Draw an arc of the earth; Draw a radius from the center then thru the arc
to a point 380 miles above the arc.
Draw a tangent from the top of the previous segment to a point that
is tangent to the arc.
Draw a radius from the center to the point of tangency (that should form
a right angle at the point of tangency).
Label the angle of depression from the spacelab 65.8 degrees.
-------------------------
EQUATION:
sin(65.8) = r / (r+380)
0.91212(r+380) = r
0.91212r + 346.6056 = r
346.6056 = 0.08788r
r = 3944.078.. miles
=========================
Cheers,
Stan H.